# Deep learning reveals key predictors of thermal conductivity in covalent organic frameworks

### v1

@@ Results
The same trend is observed for numerous examples, with additional ones presented in the Supplementary Information Section 5. We conclude that this increased attention is indicative of the deep learning model’s understanding that these dangling masses are significant in predicting thermal conductivity. Furthermore, it suggests that the presence of dangling masses disrupts the heat transfer pathways and thereby reduces the thermal conductivity through the material. A similar effect of dangling mass lowering the thermal conductivity has been reported previously in singular polymer chains [52] and amorphous polymers, [53] but remains a novel predictor for COFs which are ordered polymeric structures.
To further validate the structure-property relationship interpreted from the deep learning model, we examine the impact of dangling atoms on the vibrational density of states (VDOS) for the aforementioned contrasting COF examples. VDOS, which characterizes the distribution of vibrational modes in a system as a function of frequency, is known to impact thermal transfer properties. Overlaps in VDOS profiles between different atoms have been shown to affect these properties in both COFs [54] and MOFs. [55]
Regression model: Random Forest Gradient Boosting XGBoost AdaBoost
Without DMR as feature $R^{2}$ $\mathbf{0.592\pm 0.039}$ $0.536\pm 0.038$ $0.551\pm 0.060$ $0.312\pm 0.103$
$R^{2}$ $\mathbf{0.728\pm 0.069}$ $0.688\pm 0.070$ $0.691\pm 0.035$ $0.416\pm 0.125$
Gini: Density $0.397$ $0.433$ $0.425$ $0.478$
Gini: DMR $0.278$ $0.286$ $0.296$ $0.243$
With DMR as feature Gini: LPD $0.185$ $0.218$ $0.146$ $0.084$
Gini: Void fraction $0.070$ $0.032$ $0.076$ $0.017$
Gini: Surface area $0.069$ $0.031$ $0.177$ $0.057$
Table 1: Comparison of performances of standard ensemble regression models at predicting $\kappa$ by excluding (top table) and including (bottom table) DMR as an input feature. We performed a 10-fold cross-validation to obtain the mean and standard deviations in performances. Gini Importance for all ensemble regression models (bottom table) for all features (including DMR) are listed below in descending order.
Through MD simulations, the VDOS is calculated using the Fourier transform of the normalized velocity autocorrelation function of specific groups of atoms (see Supplementary Information Section 6 for details). Figure 4 (third column) show the VDOS profile of carbon, nitrogen, and oxygen atoms in the representative COF examples. For any COF, we define a VDOS overlap metric $S$ as the ratio of the area under the curve (AUC) of the minimum VDOS across all atom types (at each frequency) and the AUC of the maximum VDOS across all atom types (at each frequency), i.e.,
@@
Here, $\omega$ denotes the frequency; $\mathrm{f_{C}^{V}}(\omega)$, $\mathrm{f_{N}^{V}}(\omega)$, $\mathrm{f_{O}^{V}}(\omega)$, and $\mathrm{f_{B}^{V}}(\omega)$ denote the VDOS at frequency $\omega$ for the carbon, nitrogen, oxygen and boron atoms, respectively. Additionally, we differentiate between atoms in the main branch and dangling atoms, where the VDOS profiles for dangling atoms are denoted with a (.$(.)^{\text{d}}$ superscript. A large overlap in VDOS profiles of different atom types would result in $S\approx 1$, whereas a small overlap would yield $S\approx 0$. A high overlap indicates that phonon waves (i.e., quantized modes of vibrations responsible for thermal energy transfer within the crystal lattice) can propagate freely throughout the structure. This would lead to minimized phonon scattering and facilitate efficient heat transfer across the material.[56]
In the COF structures, which lack or have minimal dangling atoms, the VDOS profiles of the atoms are broad and overlap significantly, e.g., $S\approx 0.41$ and $0.49$ for examples in Figure 4a and c, respectively. This results in an even distribution of vibrational modes across a wide frequency spectrum for all atoms. In contrast, the COFs with substantial dangling atoms exhibit a VDOS profile with minimal overlap, e.g., $S\approx 0.08$ and $0.09$ for examples in Figure 4b and d, respectively. The vibrational modes of individual atoms in these COFs are confined to much narrower frequency bands, which leads to a less harmonized VDOS profile. The lack of overlap suggests that the vibrations are highly localized around the dangling atoms, which causes the phonon waves to be scattered and disrupting their ability to transfer heat efficiently through the material.[57] We interpret the mismatch in the VDOS profiles as an energy barrier for phonon transport, where phonons encounter resistance, preventing smooth energy transfer across atoms. This suggests that the absence of dangling mass is crucial for enhancing thermal conductivity in COFs. Dangling atoms create these mismatches in VDOS, which in turn hinders phonon transport and reduces thermal conductivity.
@@
We then introduce DMR as an additional feature, alongside pore size, density, void fraction, and surface area, into the classical ensemble regression models that served as the baseline for PMTransformer. The inclusion of DMR results in a marked improvement in the goodness-of-fit (see Table 1) across all the models, i.e., random forest, gradient boosting, XGBoost and AdaBoost. To assess the importance of individual features, including DMR, we analyzed their Gini importance.[58] These values quantify the contribution of each feature to the overall predictive power of the model by measuring how much each feature reduces the Gini impurity or randomness when making decisions in the ensemble model. We note that DMR emerges as the second most influential predictor with 24-30% importance, closely following behind density (see Table 1). This underscores the contribution of the DMR to the models’ predictive capabilities and the role of dangling atoms as a novel predictor for thermal conductivity in COFs.
@@
The screening is performed with the objective of facilitating the identification COFs with extremely high or low thermal conductivity without resorting to computationally expensive simulations. We observe an average screening rate of 0.07 seconds per COF structure (see Supplementary Information Section 7 for details on computational runtimes), thereby enabling a speedup of almost seven orders of magnitude compared to MD-based screening.
@@
Once potential candidates with extreme thermal conductivities are identified using the PMTransformer, we then calculate their thermal conductivities through MD simulations for these selected few, rather than for the entire datasets. The results are summarized in Figure 5. Our screening process on the dataset by Mercado et al.[41] revealed COFs exhibiting thermal conductivities up to twice the maximum value observed in the training set, but stemming from the same design space. By screening the dataset by Lan et al.,[59] we discover COFs with thermal conductivities five times as high as the ones in our training set. Additionally, we discovered several COFs with thermal conductivities lower than any values previously recorded in our dataset. A visual inspection reveals that these COFs feature a combination of large pore sizes, low density, and a substantial presence of dangling masses. For all the identified COFs, we observe that thermal conductivity shows a strong trend with both dangling mass ratio (DMR) and VDOS overlap ratio ($S$), which is in agreement with the previous observations.
Conclusion
@@
In this study, we compiled an extensive dataset of thermal conductivities for 2,471 two-dimensional COFs, computed using NEMD. Despite observing some structure-property trends with conventional descriptors such as density, pore size, void fraction, and surface area, no single descriptor or combination thereof consistently predicted thermal conductivity with high accuracy, highlighting the complexity of COF structure-property relationships. To enhance prediction accuracy, we trained a transformer-based deep learning model incorporating a multi-head attention mechanism. This model significantly outperformed traditional ensemble regression models, achieving an $\mathrm{R^{2}}$ of $0.908\pm 0.007$.
@@
Further analysis using the transformer’s attention mechanism revealed that COFs with higher amounts of dangling atoms exhibited lower thermal conductivities due to disrupted heat transfer pathways, a novel and significant predictor of thermal conductivity not previously reported. Additionally, incorporating the dangling mass ratio, i.e., DMR as an extra descriptor in traditional ensemble regression models significantly improved prediction accuracy. Across different models, DMR exhibited 24-30% in Gini importance, underscoring the importance of dangling mass in structure-property relationships. The vibrational density of states analysis corroborated these findings, showing that dangling masses create mismatched vibrational modes that hinder thermal transfer.
@@
Data availability
Author contributions
P.T. and Y.Z. generated the dataset. Y.Z. developed the NEMD pipeline to compute the thermal conductivities. Y.G. developed the dangling mass computation algorithm. P.T. and Y.Z. analyzed the data. P.T. trained the PMTransformer model and analyzed the results. P.T., Y.Z., and Y.G. visualized the results. S.K. and A.V. conceived and supervised the research. All authors contributed to writing the manuscript and approved the final version.
Competing interests
The authors declare no competing interests.
Acknowledgements
P.T., Y.G., and S.K. acknowledge that this material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7020. P.T., Y.G., and S.K. acknowledge that this work was supported in part by Oracle Cloud credits and related resources provided by Oracle for Research. A.V. and Y.Z. acknowledge Microsoft Climate Research Initiative and University of Washington for funding. We thank Prof. Marcel Sluiter of TU Delft for his helpful discussions and insights during the course of this research. Additionally, we thank Eric Vo of University of Washington for his help in preparing simulations of COF structures.

### v2

@@ Results
The same trend is observed for numerous examples, with additional ones presented in the Supporting Information Section 5. We conclude that this increased attention is indicative of the deep learning model’s understanding that these dangling masses are significant in predicting thermal conductivity. Furthermore, it suggests that the presence of dangling masses disrupts the heat transfer pathways and thereby reduces the thermal conductivity through the material. A similar effect of dangling mass lowering the thermal conductivity has been reported previously in singular polymer chains [54] and amorphous polymers [55], but remains a novel predictor for COFs which are ordered polymeric structures.
To further validate the structure-property relationship interpreted from the deep learning model, we examine the impact of dangling atoms on the vibrational density of states (VDOS) for the aforementioned contrasting COF examples. VDOS, which characterizes the distribution of vibrational modes in a system as a function of frequency, is known to impact thermal transfer properties. Overlaps in VDOS profiles between different atoms have been shown to affect these properties in both COFs [56] and MOFs [57].
Through MD simulations, the VDOS is calculated using the Fourier transform of the normalized velocity autocorrelation function of specific groups of atoms (see Supporting Information Section 6 for details). Figure 4 (third column) show the VDOS profile of carbon, nitrogen, and oxygen atoms in the representative COF examples. For any COF, we define a VDOS overlap metric $S$ as the ratio of the area under the curve (AUC) of the minimum VDOS across all atom types (at each frequency) and the AUC of the maximum VDOS across all atom types (at each frequency), i.e.,
@@
Here, $\omega$ denotes the frequency; $\mathrm{f_{C}^{V}}(\omega)$, $\mathrm{f_{N}^{V}}(\omega)$, $\mathrm{f_{O}^{V}}(\omega)$, and $\mathrm{f_{B}^{V}}(\omega)$ denote the VDOS at frequency $\omega$ for the carbon, nitrogen, oxygen and boron atoms, respectively. Additionally, we differentiate between atoms in the main branch and dangling atoms, where the VDOS profiles for dangling atoms are denoted with a (.$(.)^{\text{d}}$ superscript. A large overlap in VDOS profiles of different atom types would result in $S\approx 1$, whereas a small overlap would yield $S\approx 0$. A high overlap indicates that phonon waves (i.e., quantized modes of vibrations responsible for thermal energy transfer within the crystal lattice) can propagate freely throughout the structure. This would lead to minimized phonon scattering and facilitate efficient heat transfer across the material [58].
In the COF structures, which lack or have minimal dangling atoms, the VDOS profiles of the atoms are broad and overlap significantly, e.g., $S\approx 0.41$ and $0.49$ for examples in Figure 4a and c, respectively. This results in an even distribution of vibrational modes across a wide frequency spectrum for all atoms. In contrast, the COFs with substantial dangling atoms exhibit a VDOS profile with minimal overlap, e.g., $S\approx 0.08$ and $0.09$ for examples in Figure 4b and d, respectively. The vibrational modes of individual atoms in these COFs are confined to much narrower frequency bands, which leads to a less harmonized VDOS profile. The lack of overlap suggests that the vibrations are highly localized around the dangling atoms, which causes the phonon waves to be scattered and disrupting their ability to transfer heat efficiently through the material [59]. We interpret the mismatch in the VDOS profiles as an energy barrier for phonon transport, where phonons encounter resistance, preventing smooth energy transfer across atoms. This suggests that the absence of dangling mass is crucial for enhancing thermal conductivity in COFs. Dangling atoms create these mismatches in VDOS, which in turn hinders phonon transport and reduces thermal conductivity.
To further confirm the effect of dangling atoms on phonon dispersion in COFs, we calculate the phonon spectral energy density (pSED) of the example COFs and their dangling counterparts, as presented in Section 5 in Supporting Information. The pSED of COFs containing more dangling atoms exhibits higher magnitudes and broadening bands, indicating stronger anharmonicity and vibrational scattering in the vibrational modes, consequently resulting in reduced thermal conductivity [60].
@@
We then introduce DMR as an additional feature, alongside pore size, density, void fraction, and surface area, and train classical ensemble regression models, such as the Random Forest [61], Gradient Boosting [62], XGBoost [63], and AdaBoost [64] and outline their performances in Table 2. To assess the importance of individual features, including DMR, we analyzed their Gini importance and conducted a permutation feature importance analysis [65]. These values quantify the contribution of each feature to the overall predictive power of the model. The Gini importance measures how much each feature reduces the Gini impurity or randomness when making predictions in the ensemble model [61]. On the other hand, permutation feature importance measures the impact of a feature by assessing how much the model’s performance decreases when the values of that feature are randomly shuffled. A greater drop in performance indicates that the feature is more important in predicting the target variable [61]. We note that for the best performing regressor, the Random Forest, DMR emerges as the second most influential predictor with 27.8% Gini importance, closely following behind density (see Table 2). Similarly, the permutation feature importance analysis also highlights DMR as a key feature, further supporting its significant role in predicting thermal conductivity, with its importance consistently ranking just below density. Finally, we conducted SHAP analysis [66] on the the Random Forest to better understand the impact of each feature on the individual test data points. As shown in Figure 5, the analysis reveals that higher DMR values are associated with lower predicted $\kappa$, confirming the trend we observed. The consistency of these findings across different methods—Gini, permutation feature importance, and SHAP—reinforces the importance of DMR as a key factor in predicting thermal conductivity.
Regression Model $R^{2}$
Random Forest $\mathbf{0.728\pm 0.069}$
Gradient Boosting $0.688\pm 0.070$
XGBoost $0.691\pm 0.035$
AdaBoost $0.416\pm 0.125$
Table 1: Performances of standard ensemble regression models at predicting $\kappa$ by including DMR as an input feature. The table reports $R^{2}$ scores obtained through 10-fold cross-validation.
Feature importances for Random Forest
Feature Gini PFI
Density 0.397 0.670
DMR 0.278 0.511
LPD 0.185 0.455
Void Fraction 0.070 0.093
Surface Area 0.069 0.022
Table 2: Feature importances for the best-performing regression model, i.e., Random Forest. The table reports both Gini and Permutation Feature Importance (PFI) values for all features.
Figure 5: Distribution of the effect of each feature on the model’s predictions. Each point represents a SHAP value for a feature across all samples, with the color intensity indicating feature values (ranging between minimum and maximum values). The position along the x-axis represents the magnitude of the feature’s impact.
@@
The screening is performed with the objective of facilitating the identification COFs with extremely high or low thermal conductivity without resorting to computationally expensive simulations. We observe an average screening rate of 0.07 seconds per COF structure (see Supporting Information Section 7 for details on computational runtimes), thereby enabling a speedup of almost seven orders of magnitude compared to MD-based screening.
@@
Once potential candidates with extreme thermal conductivities are identified using the PMTransformer, we then calculate their thermal conductivities through MD simulations for these selected few, rather than for the entire datasets. The results are summarized in Figure 6. Our screening process on the dataset by Mercado et al. [47] revealed COFs exhibiting thermal conductivities up to twice the maximum value observed in the training set, but stemming from the same design space. By screening the dataset by Lan et al. [67], we discover COFs with thermal conductivities five times as high as the ones in our training set. Additionally, we discovered several COFs with thermal conductivities lower than any values previously recorded in our dataset. A visual inspection reveals that these COFs feature a combination of large pore sizes, low density, and a substantial presence of dangling masses. For all the identified COFs, we observe that thermal conductivity shows a strong trend with both dangling mass ratio (DMR) and VDOS overlap ratio ($S$), which is in agreement with the previous observations.
Discussion
@@
In this study, we compiled an extensive dataset of thermal conductivities for 2,471 two-dimensional COFs, computed using NEMD. Despite observing some structure-property trends with conventional descriptors such as density, pore size, void fraction, and surface area, no single descriptor or combination thereof consistently predicted thermal conductivity with high accuracy, highlighting the complexity of COF structure-property relationships. To enhance prediction accuracy, we trained a transformer-based deep learning model incorporating a multi-head attention mechanism. This model significantly outperformed traditional ensemble regression models, achieving an $\mathrm{R^{2}}$ of $0.909\pm 0.006$.
@@
Further analysis using the transformer’s attention mechanism revealed that COFs with higher amounts of dangling atoms exhibited lower thermal conductivities due to disrupted heat transfer pathways, identifying a novel and significant predictor of thermal conductivity. The random forest regressor, identified as the best-performing regression ensemble model, was analyzed using Gini importance, permutation feature importance, and SHAP values. These analyses consistently highlighted the dangling mass ratio as the second most important predictor of thermal conductivity. The vibrational density of states analysis further supported these findings, showing that dangling masses introduce mismatched vibrational modes that hinder thermal transfer.
@@
Methods
The molecular dynamics setup, the distribution of in-plane thermal conductivities, and feature correlations (including DMR) (Section 1); implementation details of the ensemble regression models and regression performances excluding DMR (Section 2); the deep learning setup (Section 3); the dangling mass computation (Section 4); additional comparisons of COFs with and without dangling mass, including phonon dispersion maps for all examples (Section 5); details on the VDOS and pSED computation (Section 6); and estimated runtimes (Section 7) are summarized in the Supporting Information.
Acknowledgements
P.T., Y.G., and S.K. acknowledge that this material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7020. P.T., Y.G., and S.K. acknowledge that this work was supported in part by Oracle Cloud credits and related resources provided by Oracle for Research. A.V. and Y.Z. acknowledge Microsoft Climate Research Initiative and University of Washington for funding. The funders played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
We thank Prof. Marcel Sluiter of TU Delft for his helpful discussions and insights during the course of this research. Additionally, we thank Eric Vo of University of Washington for his help in preparing simulations of COF structures.
Competing interests
The authors declare no competing interests.
Data availability